OpenStudy (anonymous):

Please help me, I don't know where I went wrong and I keep getting this same answer! *WILL FAN AND MEDAL*! Please help me check my answer, PLEASE!

OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (-4, 0) and a directrix at x = 4. A. y^2 = -8x B. 16y = x^2 C. y = -1/16x^2 D. x = -1/16y^2

OpenStudy (anonymous):

I keep getting y^2 = -1/16x

OpenStudy (anonymous):

or x = -16y^2

OpenStudy (anonymous):

@misty1212 @satellite73 @Peaches15 anybody??

OpenStudy (anonymous):

b

OpenStudy (anonymous):

How do you know? @kali_sky

OpenStudy (anonymous):

We know that the directrix is x=4 so the equation must be y^2, and we know that the focus is (-4,0) so the center of the parabola is (0,0) and p=-4, then using the formula \[\Large (y-k)^{2}=4p(x-h)\] of the parabola with center (h,k) and p=semi-axis length, replacing we get \[\Large(y-0)^{2}=-4\times4(x-0)\] \[\Large y ^{2}=-16x\]